An aspirin product on the market claims to use magnesium hydroxide, Mg(OH)2, as a buffering agent. How could you verify whether magnesium ions are present?
1
Expert's answer
2012-01-27T08:51:22-0500
1. You need add some acid (HCl) to aspirinem(if Mg(OH)2 is present it must react with acid)
Mg(OH)2 + 2HCl = MgCl2 + 2H2O
2. Than we need add Na2HPO4
MgCl2 + Na2HPO4 = 2NaCl + MgHPO4 (it precipitates)
precipitate should be white and crystalline

Also Mg2+ reacted with Titanium yellow, C28H19N5Na2O6S4,giving red precipitate.